Shift report 20000903 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Sep 03 2000 - 00:14:31 EDT

  • Next message: Michael Murray Tel 409 845 1411,FAX 1899, Home 823 3386: "BB ADC gain and multiplicity calculation"

    ------------------------------------------------------------
    Date: Sun Sep  3 00:14:31 2000, Shift: 16:00-24:00
    Supervisor: JHL, with: AJ
    Log book#5 pages 77 - 80
    
    Shift Summary:
    ------------------------------------------------------------
    Rainy Saturday.
    
    1.  Beam
    Machine recovered from the power dip. 
    Stable beam since 20:00 at blue=14*10^12 yellow=15*10^12.
    
    2.  Data
    Run 2513 (20:00-22:00) has ~ 100K collisions.
    Run 2514 (22:00-24:00) has ~  90K collisions.
     
    3.  Setting and Trigger
    MRS: 40 degree B pol
    FS: 4 degree B pol
    Trigger: 1,4,5,6,8 (sd: 2,200,5000,1,1) 
    (changed scale down for trigger 4: 40->200
    trigger ratio trigger4:trigger6=1:0.8)
    
    4.  Detector Problems
    TOFW HV mainframe MRS1440 went off (due to power the dip?)
    brought back up except 3 channels (1,5,7): notified K.O.
     
      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Sep 03 2000 - 00:14:53 EDT